OpenStudy (anonymous):

What graph would match this : x – 3y = –12 2x – y = 1 System of Equations

OpenStudy (anonymous):

Do you need to solve the system?

OpenStudy (anonymous):

IWould you like me to describe the graph or...?

OpenStudy (anonymous):

Describe the graph

OpenStudy (anonymous):

Oh ok one sec

OpenStudy (anonymous):

one line goes from (0,4) to (-12,0)

OpenStudy (anonymous):

The other goes from (0,1) to (1,1)

OpenStudy (anonymous):

Can you tell me how you solved that ?

OpenStudy (anonymous):

Sure \[x-3y=-12 \rightarrow -3y=-x-12 \rightarrow (divide, by, negative, three) y=1/3x+4 \]

OpenStudy (anonymous):

thats the first equation all you do is isolate y and put it in slope intercept form

OpenStudy (anonymous):

Then from there you graph the y-intercept and use the slop to find the next point

OpenStudy (anonymous):

Ok so would this be the graph

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np